OpenStudy (anonymous):

Determine the factors of x2 − 6x − 8. Prime (x + 4)(x − 2) (x − 4)(x + 2) (x + 8)(x − 1)

OpenStudy (zzr0ck3r):

foil them

OpenStudy (saifoo.khan):

How to FOIL the first option? :D

OpenStudy (anonymous):

\[x^2 − 6x − 8\] what multiplies to -8 but adds to -6

OpenStudy (anonymous):

Multiplies to -8 : -1 * 8 = -8 -2 * 4 = -8 1 * -8 = -8 2 * -4 = -8 Adds to -6 : -1 + 8 = 7 -2 + 4 = 2 1 + -8 = -7 2 + -4 = -2

OpenStudy (anonymous):

You could check the discriminant. If b^2 - 4ac isn't a perfect square, you won't get rational roots, i.e., it will be prime.

OpenStudy (anonymous):

In this case, the discriminant seems to be about 68, which would indicate irrational roots. The quadratic is prime.

OpenStudy (anonymous):

The above generalizations will hold true for integer coefficients.

OpenStudy (anonymous):

@jen1511 do you understand?